### v2

# Mind Your Theory: Theory of Mind Goes Deeper Than Reasoning

Abstract
Theory of Mind (ToM) capabilities in LLMs have recently become a central object of investigation, sparking debates and discussions. In this position paper, we explore many lines of work in different communities in AI and cognitive science. Inspired by cognitive work, we view ToM tasks as a two-step process: (I) first, determining whether and how to invoke ToM, which includes setting the appropriate Depth of Mentalizing (DoM); and (II) second, applying correct inference given the appropriate DoM. We identify that many works about ToM in LLMs, such as benchmarks and add-on modules, tend to unjustly overlook the first step and focus exclusively on the second one, which can be framed as static logic problems. We support our empirical evidence about the difficulty of the different steps in existing benchmarks. We conclude with suggestions for improved evaluation of ToM capabilities, inspired by dynamic environments used in cognitive tasks in biological agents.
Mind Your Theory: Theory of Mind Goes Deeper Than Reasoning
Eitan Wagner1  Nitay Alon1,2  Joseph M. Barnby3  Omri Abend1 1 Hebrew University of Jerusalem  2 MPI for Biological Cybernetics 3 Royal Holloway University of London {eitan.wagner, nitay.alon}@mail.huji.ac.il
**footnotetext: Equal contribution.
1 Introduction
The ability of Large Language Models (LLMs) to efficiently integrate social information is essential for ensuring AI trust and safety, as well as reasoning about situations that involve multiple agents. This capability relies on inferring and representing other synthetic and biological agents’ beliefs, desires, and intentions, also known as mentalizing, or Theory of Mind (ToM) Premack and Woodruff (1978).
LLMs have been recently regarded as general-purpose reasoning models Brown et al. (2020), and their ToM capabilities have come under scrutiny, particularly as to whether social cognition can emerge purely from associative principles Sap et al. (2022). Notably, Kosinski (2023) claimed that ToM capabilities emerge in post-GPT3 models, like ToM in children Astington and Jenkins (1995). This claim sparked debates on the correct interpretation of these results Ullman (2023); Pi et al. (2024).
Figure 1: Overview of ToM error types. Existing benchmarks address cases of clear ToM intention and deal mainly with step II (and Type C errors), overlooking the possibility of Type B errors. In contrast, the decision to invoke ToM (step I) should address Type A and Type B errors.
Existing benchmarks for ToM primarily concentrate on whether agents have correct beliefs about others Kosinski (2024). However, within the context of social interaction, ToM capabilities involve two steps Schaafsma et al. (2015); Leslie et al. (2004): (1) determining the depth of mentalization (DOM; Barnby et al., 2023) to use, which specifically includes the decision of whether to model the agents separately (self-other divide); and (2) applying correct inference for each agent’s mental state, given the depth. Failure in any step leads to an incorrect conclusion (see Figure 1). We argue that very little attention is paid to whether LLMs correctly invoke ToM (when needed). Determining this is critical: (1) it ignores a critical theoretical component of ToM that has been well-defined in humans; (2) it confounds ToM errors with poor or fuzzy logic Pi et al. (2024), and (3) the act of mentalizing has a price, both in resources and performance Rilling et al. (2004); Keysar et al. (2003); Devaine et al. (2014), and so more precisely understanding moment-to-moment ToM in LLMs may improve compute efficiency at equivalent parameter sizes. Of note, we do not focus on the evolutionary pressures and resource-rational necessities of social representation – the why of ToM – and merely focus on the how.
In this paper, we first discuss contemporary approaches to ToM in AI research at a high level. This includes LLM benchmarks, ToM add-ons for LLMs, and agent-focused ToM models. We identify that the focal point of current LLM and AI research is logical inference Verbrugge and Mol (2008), corresponding to step II (and Type C errors), which is based on the determination of DoM in the first stage. We experiment with two common ToM benchmarks to demonstrate this conclusion. We then turn to the cognitive science literature and identify differences in the definition of ToM behind biological and synthetic agents, as the former focuses also on step I – the act of mentalizing, and not only on correct inference. We finally identify ways to improve ToM research in AI, inspired by work in biological agents.
Our contributions are: (1) We explore and discuss ways of addressing ToM in various fields. (2) We formally describe ToM tasks as a two-step task, inspired by cognitive science. (3) We identify a lacuna in existing LLM works which tend to overlook the first step of mentalizing and focus instead on logic problems. (4) We provide a roadmap for appropriate evaluation of ToM invocation inspired by biological agents.
2 ToM in AI research
Here we analyze the various approaches and challenges used to model and assess ToM in the AI communities. We begin with a review of existing AI methods before discussing LLM-focused approaches.
2.1 Formal Architectures of ToM
Inspired by Reinforcement Learning (RL), ToM has been modeled as an inferential process in which an agent makes inferences about other agents Baker et al. (2011). This approach is similar to Inverse-RL Ng and Russell (2000); Jara-Ettinger (2019), where an inferring agent learns an association between actions and unobserved variables, such as goals or future actions.
The main characteristic of ToM is the depth of recursion, also known as Depth of Mentalizing (DoM). In observational tasks (e.g., SA tasks), the inferring agent is typically modeled as level-zero Depth of Mentalizing (DoM) – DoM$(0)$ Barnby et al. (2023). This model does not allow for recursive beliefs (“I think that you think that I think”). Instead, it is limited to shallower mentalizing (“I think that”) - with an absence of expected knowledge of the observer (see Figure 2).
Higher DoM levels (like the one presented in Wu et al. (2023)) are modeled using Interactive-POMDPs (IPOMDP; Gmytrasiewicz and Doshi, 2005), which model other agents as part of the world model. We refer the reader to Alon et al. (2023) for a description of the model and its applications in multi-agent environments.
Some attempts have been made to integrate recursive formulations into LLMs. However, due to the black-box nature of LLMs, this is difficult to achieve Zhang et al. (2023); Hu and Shu (2023).
2.2 ToM Benchmarking
Evaluation of LLMs on ToM benchmarks is an active line of research. A dominant component in these tests is variations of the Sally-Anne (SA) test (Wimmer and Perner, 1983; Baron-Cohen et al., 1985). In the SA test, participants are tested on their ability to identify false beliefs, by recognizing that Anne has moved a ball while Sally, who is not exposed to this event, maintains the belief that the ball did not move. These correspond to observational inference tasks Miura and Zilberstein (2021).
Variations in the SA task have been used as a gold standard benchmark for ToM in LLMs Grant et al. (2017); Nematzadeh et al. (2018); Gandhi et al. (2023). Benchmarks like ToMi Le et al. (2019) and OpenToM Xu et al. (2024), improve previous benchmarks with less predictable generation protocols and more natural settings. Rather than vignette-based tasks, some benchmarks use a language-based conversational setting to assess ToM reasoning, such as Mindcraft Bara et al. (2021), FANToM Kim et al. (2023), and NegotiationToM Chan et al. (2024). Some benchmarks extend the SA task to more complex objectives. This includes higher-order ToM, as in HiToM Wu et al. (2023); additional perception inference, as in Percept-ToMi and Percept-FANToM Jung et al. (2024); and epistemic logic conclusions Sileo and Lernould (2023). Additional extensions derive questions from common ground annotations, as in CommonToM (Soubki et al., 2024) or from other modalities, as in MMToM-QA (Jin et al., 2024). lity of these tasks in assessing ToM in LLMs is still a matter of debate. GPT3 showed difficulties engaging with ToMi Sap et al. (2022), whereas more complex models have been argued to show more success Kosinski (2023); Gandhi et al. (2023). Nevertheless, slight alterations to task structure can alter the performance Ullman (2023); Shapira et al. (2024); Ma et al. (2023a); Chen et al. (2024); Nickel et al. (2024). The theoretical implications of these effects are therefore still under debate Pi et al. (2024). Other criticisms of current benchmarks are the lack of transparency in the training set used to compare LLMs with human-like performance Ma et al. (2023b); Shapira et al. (2024).
In particular, all these benchmarks are static in the sense that they deal with pure observation rather than the evolution of the moment-to-moment model during interaction Ma et al. (2023b). They therefore only distantly resemble the application of ToM capabilities in real-world settings. Additionally, these benchmarks assume that there is always a single correct answer, rather than a range of possible beliefs and behaviors.
2.3 ToM Add-ons
Many methods were proposed for improving ToM capabilities in existing architectures. We describe these methods as ToM Add-ons. These components attempt to complement LLMs with external modules, assuming it is necessary.
There are various types of suggested add-ons. Appropriate prompting techniques have shown improved scores on classic ToM tasks Tan et al. (2024); Jung et al. (2024); Huang et al. (2024); Wilf et al. (2024). Explicit symbolic modules can also be updated based on input and improve accuracy Qiu et al. (2024); Hou et al. (2024). Others add post-model decoding methods Sclar et al. (2023) or symbolic planners Jin et al. (2024) to improve performance.
Figure 2: Hierarchical Theory of Mind (ToM). Depth of Mentalizing (DoM) indicates how much an agent (blue) recursively nests beliefs about their partner (orange). At DoM(-1) the agent only considers their own policy ($\pi$). DoM(0) beliefs consider $\pi$ and a partner’s $\pi$. DoM(1) beliefs consider DoM(0) and the beliefs of the partner about the agent’s $\pi$.
3 What Can We Learn From biological Agents?
Cognitive science, Economics, and Neuroscience have explored ToM in humans for decades. Insights gained over the years can inform the development of architectures in LLMs and AI in general, but regrettably, this literature is seldom referenced in NLP. Particularly, in these fields researchers explored adaptability — changes in the DoM level of an agent during an interaction in response to feedback (Grosskopf and Nagel, 2007) and intentionality — invoking ToM when needed and avoiding it when unneeded (Aichhorn et al., 2006).
[…]
3.1 Invoking DoM
Several theoretical and empirical works identify the common principles of social interaction. Biological agents performing cooperative tasks can afford very shallow mentalizing Devaine et al. (2014), and instead rely on a common policy (e.g., social norms or prosocial goals). When competition between agents is present, hierarchical mentalizing is beneficial to avoid duplicitous action Alon et al. (2023). However, moving from a shallow to a hierarchical level of mentalizing is more expensive, requiring additional time and energy to compute (Sweller, 1988; Bossaerts and Murawski, 2017). This limits the ability to perform ToM-demanding tasks for a long time (Rilling et al., 2004), and can lead to failures in social reasoning when ToM is needed but jeopardized by disconnected typical neural circuits Barnby et al. (2022). While LLMs do not have the same energy constraints as humans, failures to appropriately capture the degree to which LLMs can move into competitive modes of mentalizing may hinder human-AI alignment.
Even with properly invoked ToM reasoning, an agent may still suffer from maladaptive DoM. For example, remaining fixed in a shallow DoM may miss malicious intent from other agents Sarkadi et al. (2019). On the other hand, invoking hierarchical DoM during a cooperative problem can induce false beliefs of harm Alon et al. (2024). Moreover, as described above, when interacting with agents with varying DoM levels (Camerer et al., 2003), a “correct” DoM level may be ill-defined.
3.2 Revisiting Existing Work
What do the ToM benchmarks benchmark?
With current vignette-based reasoning tasks, the necessity of ToM is made obvious. Thus, current benchmarks practically address step II of our formulation, thus precluding the need to use ToM or even reason about self vs. other distinctions. Even as challenging benchmarks are introduced, the focus is still on non-interactive presentations with templatized questions, in which the exact required DoM level, $k$, is rather obvious from the question (see Appendix A for a case study on two challenging benchmarks). The problems these benchmarks address are therefore logical ones, which, while challenging, ignore crucial aspects of ToM.
We note that even if $DoM$ is not given, as long as we test the prediction for $Y|X$, we cannot determine whether performance reflects ToM capabilities. If the model is incorrect, if we do not differentiate between the steps, we will not know if the failure was due to undermentalization (Type B error) or due to failed reasoning (Type C error). If the model is correct but ToM is unnecessary, we cannot tell if the success was based on the correct ToM (Type A error) Kim et al. (2023). Only if ToM is necessary and the model succeeded, we may conclude that ToM was used and correctly so (with the caveat of spurious correlations).
What do ToM add-ons add?
Typical ToM add-on works propose methods for specific ToM tasks. Essentially, these methods assume ToM is required ($DoM(k\geq 0)$), thus focusing on step II. The decision about the required DoM level (step I) is not directly addressed.
What do formal models model?
Typical works on ToM propose models that inherently assume a fixed DoM$(k),k\geq 0$. While these models are beneficial for learning at this level, the fixed, immutable DoM level assumption (with some exceptions like Camerer et al. (2003) and Grosskopf and Nagel (2007)) limits the ability to properly model human data, where DoM levels may change. Moreover, they do not address the step of invoking ToM (step I) – they assume that ToM is needed and invoked for inference within a given task. A more complete model that examines both steps is, to the best of our knowledge, missing from this literature.
4 Discussion and Conclusion
In this work, we cast ToM in LLMs as an interdisciplinary topic. We argue that problems that existing NLP benchmarks and AI works typically address are non-interactive logical problems. While many of the benchmarks are valuable as reasoning tests and may provide insights into the performance of LLMs, they do not address the core issue of ToM in LLMs: interaction, and social adaptation.
We argue that a better understanding of ToM in LLMs requires: (1) benchmarks that test whether DoM is correctly invoked or not, (2) evaluation for ToM in interactive settings where the model is an active agent, and (3) evaluation for the appropriate degree of DoM, given a social context (cooperative vs competitive).
Some works introduce ToM tasks for LLMs or formal models in interactive domains Sclar et al. (2022); Li et al. (2023); Ying et al. (2025), albeit with simplified contexts. We believe that a meaningful understanding of ToM capabilities should be through works like these with additional domains and contexts. These tests should cover a variety of ToM complexities (e.g., with decisions about the appropriate DoM), social contexts (e.g., with varying levels of cooperation), and ToM behavior of other agents (e.g., heterogeneous DoM levels in the population).
Limitations
Our work focuses on attributed beliefs, which are a single component of ToM. Previous work has pointed out that true evaluation of ToM must include many other factors, such as emotions and desires Ma et al. (2023b). Despite this limitation, we note that belief attribution remains a critical component in ToM evaluation.
We also note that the ToM literature in cognitive science is vast and often conflicting. The components of ToM are actively debated and thus our recommendations must be interpreted in light of this.
Acknowledgments
This work was supported in part by the Israel Science Foundation (grant no. 2424/21)
Appendix A A Case Study on Existing Benchmarks
Here we empirically demonstrate our arguments on two benchmarks, HiToM Wu et al. (2023) and FANToM Kim et al. (2023). These benchmarks were shown to be challenging even for strong models. We argue that the main challenge in these benchmarks is due to the complex reasoning required (type C errors) and not due to ToM.
To test this, we design a simple test to predict the order of ToM in a question. In both benchmarks, samples are labeled with the order of ToM required to answer the question. If models can determine the correct order of ToM but are incapable of answering the question, this indicates that the failure is not in ToM invocation.
A.1 Experimental Setup
Data.
HiToM is a benchmark that is based on Sally-Anne-like stories. The stories include higher-order beliefs and questions about the stories range from factual questions (order 00) to questions about $4$-th order beliefs. Each story has many questions in many orders. Additionally, HiToM has stories with communication and deception, such that in many cases story characters intentionally report false statements.
FANToM consists of stories with conversations in which characters enter and leave. The questions are various. They include factual (order 00) and belief questions, with belief questions including first- and second-order beliefs (the data includes the question type).
For evaluation, we use a set of $100$ random examples from each dataset. Some manual experiments were done, using a small number of examples. These examples were excluded from the evaluation.
Settings.
We test whether the models can predict the ToM level of a question. We test once with the question alone and once with the story too. For each of these tests, we test with or without Chain of Thought (CoT) in the response.
Models.
We test three models: GPT-3.5-turbo, GPT-4o-mini, and GPT-4o. [^1] https://platform.openai.com/docs/models/models
For HiToM, Wu et al. (2023) report an overall accuracy of $0.315$ with GPT-3.5-turbo and $0.59$ with GPT-4. For FANToM, an accuracy of $0.101$ is reported with GPT-3.5-turbo, $0.426$ with GPT-4-turbo, and $0.497$ with GPT-4-turbo. [^2] https://github.com/skywalker023/fantom These FANToM results are for the belief task with a short version of the story.
Benchmark Model Without Story With Story
No-CoT With-CoT No-CoT With-CoT
GPT-3.5-turbo 0.43 0.45 0.43 0.45
HiToM GPT-4o-mini 0.81 0.81 0.81 0.81
GPT-4o 0.92 1. 0.92 1.
GPT-3.5-turbo 0.52 0.53 0.52 0.53
FANToM GPT-4o-mini 0.84 0.84 0.84 0.84
GPT-4o 0.86 0.92 0.86 0.92
Table 1: ToM degree prediction results. We report the accuracy in multiple settings over $100$ random samples.
Prompt.
Here we describe the prompts used for the different settings.
For the setting with the story, we start with:
I will give you a story with a question.
For the question alone, we start with:
I will give you a question about an unseen story.
Then we give the explanation (in all cases):
I want you to tell me the Theory of Mind (ToM) order of the question. The ToM order is number of mentalization levels required to correctly answer the question. Where mentalization levels are the number of nested beliefs that one has to consider. For example, factual questions (e.g., "What day is it today?") require 0 levels of ToM, as they are invariant to beliefs. Questions about what someone else thinks (e.g., "What day does John think it is today?") require one level of ToM, as the fact itself is insufficient to answer the question. Additional levels are possible, for example, “What does John think that Mary thinks about what day it is today?” requires two levels of mentalization.
Notice that I want the ToM order required to answer the specific question and not all mentalization processes that occur. So, for example, if the story is about John thinking about something and the question is whether John is thinking, the ToM order is 0. If the question is what John thinks, the ToM order is 1. I also stress that for this task, we only address explicit beliefs and ignore implicit ones. So, for example, if the question is "Where does John take his dog on a walk?" the ToM order is 0, even though taking the dog involves an implicit mental process.
In the no-CoT setting, we add the instruction:
Your answer must be in the format: "The answer is: <answer>" Where <answer> is a number depicting the ToM order of the question. Do NOT add any additional explanations.
In the CoT setting, we add instead:
You can write explanations, but your last line must be in the format: "The answer is: <answer>" Where <answer> is a number depicting the ToM order of the question.
In the setting with the story, we input the story and the question:
The story is: <story>
and the question is: <question>
Without the story, we input only the question:
The question is: <question>
A.2 Results
We report the results in Table 1.
GPT-4o-mini achieves good results and GPT-4o is nearly perfect. GPT-3.5-turbo struggles with the task. CoT gives some improvement.
In all cases, the scores with and without the story as input are identical.
A.3 Comments and Discussion
In FANToM there are many cases of actions that include implicit ToM but are not considered as additional levels. For example, a question like “What does [A] believe about how [B] overcomes the challenge of dealing with …” is marked as first-order since it only involves one explicit belief. However, it is also reasonable to classify this question as second-order since “dealing” implies some internal intention.
Some questions in FANToM are confusing as they contradict the answer. For example, as in the previous example, a question might ask what [A] believes about [B]’s actions, but according to the story [A] has no belief about it and this is the intended answer. This is confusing since the question suggests that [A] does have a belief.
In FANToM there is also no notion of deception – all reports are assumed to be true. These assumptions are not even counted (in the dataset construction) as ToM. In a case like the previous example, if [B] said that he deals with the challenge in some way, a question about how he deals is considered zero-order. This consideration ignores an option that [B] was not telling the truth.
HiToM contains deception in some stories. However, it is always made clear to the reader what the truth is. It is also explicitly noted in the prompt that characters may not be reporting the truth.
We also note that in the HiToM paper, some analysis was done regarding the error types, with one type indicating insufficient reasoning-depth. This is similar to our experiment, but it addresses explanations given by the model in the original task. It is possible that a model can predict the correct depth but get confused in the actual reasoning process leading to an incorrect stop. In our experiment, we test the depth independently of the question itself.
In conclusion, our experiment shows that the sub-task of predicting the ToM order is much easier than the original task. Even GPT-3.5-turbo, which struggles with this task, has better scores than on the original task. Additionally, this sub-task can be performed based on the question alone.
The results indicate that the challenging part of these benchmarks is not the ToM order (=invoking ToM) but rather the inference given the order (Type C errors). Moreover, the benchmarks address the ToM level as a prediction task with a single correct solution. As we argue in the paper, this is a limited view of ToM.